# Bridging the Social & Technical Divide in Augmentative and Alternative Communication (AAC) Applications for Autistic Adults

### v1

@@ 7. Limitations & Future Work
Although we were able to collect some variation in gender and age, all of our participants were white. This is a significant limitation to our work, which we attribute to our recruitment methods. Therefore, any insight toward AAC shared in this paper may not apply to BIPOC users. This is particularly important since some (Pope et al., 2022) have identified racial disparities in access to such communication tools.
Another limitation is that our participants were only recruited online. As we did the majority of our interviews during the pandemic, there was not much of a choice. However, because of this we are missing out on a fraction of the autistic adult population that does not use the internet or social media. This could be why our participants skew on the younger side.
@@
We plan to create a prototype considering many of the features that arose from this study to see how people actually like these components in a potential AAC system and how they can best be implemented. We will also be sure to follow our suggested guidelines to determine if they indeed work and come up with new guidelines as needed.
We would also like to encourage the rest of the AAC research community to take our insights and guidelines into consideration so that we can iterate on them as a community. This needs to include support from existing and future companies that make AAC so that communication can be improved specifically for autistic adults and not as an afterthought.
@@ 8. Conclusion
Our interviews with autistic adult AAC users revealed both current societal and technical issues encountered with AAC applications. We identified what types of input & output options people prefer for their AAC, what considerations people take into account for when selecting an AAC application or switching to a different AAC application, the benefit of using AAC, issues with accessing or using AAC, social concerns, and concerns about control over their communication. We conclude by suggesting nine guidelines for natural language processing researchers, AAC application makers, and policy makers to make AAC better suited specifically for autistic adults.
Acknowledgements

### v3

@@ 7. Limitations & Future Work
Although we were able to collect some variation in gender and age, all of our participants were white. This is a significant limitation to our work, which we attribute to our recruitment methods. Therefore, while our insights might apply to BIPOC users, they would have to be validated in future studies. This is particularly important since some (Pope et al., 2022) have identified racial disparities in access to such communication tools.
Another limitation is that our participants were only recruited online. As we did the majority of our interviews during the COVID-19 global pandemic, there was not much of a choice. Because of this we are missing out on a fraction of the autistic adult population that does not use the internet or social media. This could be why our participants skewed younger.
@@
Many existing AAC applications are created for children first, leaving autistic adults to force applications to fit their needs. Although we are certainly not the first to suggest an adult-centered design McNaughton and Bryen (2007), we hope that this work is just the beginning of an emphasis on adult-first design for autistic adults and people with other developmental disabilities. Specifically, we would like to see AAC design that 1) treats users as adults or helps users be treated as adults, and 2) supports users who do not have access to an AAC specialist. Adult-centered AAC might include features such as no recording of button presses (§5.3.1,5.8.2), integrating an ”adult” vocabulary (§5.1.3,5.3.1), developing more expressive adult TTS voices (§5.2.2), or creating symbol libraries that feel more mature (5.3.1). Supporting users without SLP access does not have to be constrained to social scaffolds and could be integrated into the design of the applications themselves. For example, they could include features for making setup easier (§5.6.2) or helping people figure out what type of AAC tool would best fit their needs (§5.6.1).
@@ 8. Conclusion
Our interviews with autistic adult AAC users revealed both current societal and technical issues encountered with AAC applications. We identified what types of input & output options people prefer for their AAC, what considerations people take into account for when selecting an AAC application or switching to a different AAC application, the benefits of using AAC, issues with accessing AAC as an adult, and concerns about control over their communication. We concluded by suggesting seven novel research directions for AAC researchers to make AAC better suited specifically for autistic adults.
  1. (1)
    Conversation Planning - Referring to what the user has previously said in the conversation
  2. (2)
    Keyboard Input - Typing integration
  3. (3)
    Opt-in Functionality - Practices for making features opt-in but still informing users that they’re there and what they do
  4. (4)
    Presentation Support - Features that enable users to give long-form presentations to larger audiences
  5. (5)
    Same-app Switching - Ways of switching between input types, interfaces, or modalities within the same app
  6. (6)
    Trans* Voices - Customizable and gender-expansive synthesized voices
We would like to encourage the rest of the AAC research community to consider our insights and suggested research directions so that we can iterate on them as a community.
Acknowledgements.
@@ B.3. Privacy Table
Number of people who would use each feature. People were the most comfortable with deliberate personalization, followed by word prediction and sentence prediction. Automatic personalization was the least liked, followed by speech recognition. Logging (verbatim or topics) were both in the middle. Noteably though, nobody said they were always okay with logging verbatim.
Figure 4. Privacy Table Results – Level of comfort for use of and data storage from 7 potential features. Participants rated each feature as either something they would never use (red), something they would only use as long as the data persisted only in their current conversation (orange), something they would use with the data persisting across conversations — with the assumption that it would be deleted over time (yellow), or something they would be okay using all the time (green). The specific descriptions for each feature can be found in Appendix B.3.